# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio

### v2

@@ 2.2 Optimality Conditions via Pontryagin’s Maximum Principle
with initial conditions $X_{0}^{*}=x_{0}>0$ and $\mathbf{Y}_{0}=\mathbf{y}_{0}\in\mathbb{R}^{k}$. Recall the definitions of the volatility matrix $\bm{\sigma}(t,\mathbf{Y}_{t})$ and the covariance structure $\bm{\Omega}$ of the combined Brownian motion $\mathbf{W}_{t}=(\mathbf{W}^{X\top}_{t},\mathbf{W}^{Y\top}_{t})^{\top}$.
@@
The application of PMP involves maximizing a Hamiltonian function. To properly account for the dependence of the wealth diffusion on the control $\bm{\pi}$, the Hamiltonian should incorporate the interaction between the costate dynamics (represented by $\mathbf{Z}$ terms) and the diffusion coefficients. Following formulations in stochastic control theory (e.g., Yong & Zhou, , 2012), the appropriate Hamiltonian $\mathcal{H}$ can be expressed as:
$\begin{split}&\mathcal{H}\bigl{(}t,x,\mathbf{y},\bm{\pi},C,\lambda,\bm{\eta},\mathbf{Z}^{\lambda X},\mathbf{Z}^{\lambda Y},\mathbf{Z}^{\eta X},\mathbf{Z}^{\eta Y}\bigr{)}\\ &\quad=e^{-\delta t}\,U(C)+\lambda\Bigl{[}x\bigl{(}r(t,\mathbf{y})+\bm{\pi}^{\top}(\bm{\mu}(t,\mathbf{y})-r(t,\mathbf{y})\mathbf{1})\bigr{)}-C\Bigr{]}+\bm{\eta}^{\top}\,\bm{\mu}_{Y}(t,\mathbf{y})\\ &\quad\quad+\mathbf{Z}^{\lambda X}\bigl{(}x\,\bm{\sigma}(t,\mathbf{y})^{\top}\bm{\pi}\bigr{)}+\text{Tr}\bigl{(}\mathbf{Z}^{\eta Y\top}\bm{\sigma}_{Y}(t,\mathbf{y})\bigr{)}+\dots\end{split}$ (7)
For clarity in deriving the FOC for $\bm{\pi}$, the crucial term involving $\bm{\pi}$ beyond the drift component is the one coupling $\mathbf{Z}^{\lambda X}$ with the wealth diffusion $x\,\bm{\pi}^{\top}\bm{\sigma}(t,\mathbf{Y}_{t})$ originating from $d\mathbf{W}_{t}^{X}$. We focus on these terms for the maximization step.
@@
    $C^{1}$-regular coefficients w.r.t. $\mathbf{y}$. All coefficient functions $\bm{\mu}_{Y},\,\bm{\sigma}_{Y},\,\bm{\mu},\,\bm{\sigma}\,$ $r$ are continuously differentiable with respect to the state vector $\mathbf{y}$. Furthermore, their partial derivatives ($\partial_{\mathbf{y}}f$) are globally Lipschitz continuous in $\mathbf{y}$ and satisfy linear growth conditions. This ensures that the drift terms in the costate BSDEs, involving derivatives like $\mathcal{H}_{\mathbf{y}}$, satisfy the required regularity conditions for the existence and uniqueness of solutions (cf. Yong & Zhou, , 2012).
@@
$\qquad\qquad-\frac{V_{x\mathbf{y}}}{X_{t}V_{xx}}\,\Sigma^{-1}(t,\mathbf{Y}_{t})\,\bm{\sigma}(t,\mathbf{Y}_{t})\,\rho\,\bm{\sigma}_{Y}^{\!\top}(t,\mathbf{Y}_{t}).$
Here, $\Sigma(t,\mathbf{Y}_{t})=\bm{\sigma}(t,\mathbf{Y}_{t})\Psi\bm{\sigma}(t,\mathbf{Y}_{t})^{\top}$ represents the instantaneous covariance matrix of the risky asset returns. The value function derivatives $V_{x}=\lambda_{t}^{*}$, $V_{xx}=\partial_{x}\lambda_{t}^{*}$, and $V_{x\mathbf{y}}=\partial_{\mathbf{y}}\lambda_{t}^{*}$ are evaluated along the optimal path $(t,X_{t}^{*},\mathbf{Y}_{t})$. The formula clearly separates the portfolio into two components: the myopic demand, which optimizes the immediate risk-return trade-off as if the investment opportunity set were constant, and the intertemporal hedging demand, which adjusts the portfolio to mitigate risks arising from future stochastic changes in the state variables $\mathbf{Y}_{t}$. The hedging demand’s magnitude and direction depend on the correlation $\rho$ between asset returns and state variable innovations, and crucially on $V_{x\mathbf{y}}$, which measures how the marginal value of wealth changes with the state variables.
@@
$\pi_{t}^{*}=-\frac{V_{x}}{X_{t}V_{xx}\sigma^{2}}\bigl{(}\mu(t,Y_{t})-r(t,Y_{t})\bigr{)}-\frac{V_{xy}}{X_{t}V_{xx}\sigma^{2}}\sigma\rho\sigma_{Y},$
@@ 2.3.1 A Multi-Asset Model with One OU Factor
We first consider a model with $n$ risky assets whose risk premia are driven by a single Ornstein–Uhlenbeck (OU) factor. Specifically, each asset $i$ has a constant volatility $\sigma_{i}>0$ and a time-varying risk premium proportional to an OU state variable $Y_{t}$, such that the expected return is $r+\alpha_{i}\sigma_{i}Y_{t}$, where $r$ is the constant risk-free rate and $\alpha_{i}$ is a loading coefficient. The state variable $Y_{t}$ follows the OU process:
@@
The Brownian motions driving the assets, $\mathbf{W}_{t}^{S}=(W_{1,t}^{S},\dots,W_{n,t}^{S})^{\top}$, have an $n\times n$ covariance matrix $\Psi dt$, i.e., $\mathrm{Cov}(d\mathbf{W}_{t}^{S})=\Psi\,dt$. Furthermore, each $dW_{i,t}^{S}$ may have a correlation $\rho_{i}$ with $dW_{t}^{Y}$; let $\bm{\rho}=(\rho_{1},\dots,\rho_{n})^{\top}$.
@@
$dX_{t}=\left[X_{t}\left(\left(1-\bm{\pi}_{t}^{\top}\mathbf{1}\right)r+\bm{\pi}_{t}^{\top}(r\mathbf{1}+\bm{\alpha}\bm{\sigma}Y_{t})\right)-C_{t}\right]dt+X_{t}\bm{\pi}_{t}^{\top}\mathrm{diag}(\bm{\sigma})\,d\mathbf{W}_{t}^{S},$
where $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{n})^{\top}$, and $\bm{\sigma}=(\sigma_{1},\dots,\sigma_{n})^{\top}$. The instantaneous covariance matrix of the $n$-asset returns is $\Sigma=\mathrm{diag}(\bm{\sigma})\,\Psi\,\mathrm{diag}(\bm{\sigma})$.
@@
$\bm{\pi}_{t}^{*}=-\frac{1}{X_{t}^{*}(\partial_{x}\lambda_{t}^{*})}\Sigma^{-1}\left\{\lambda_{t}^{*}(\bm{\alpha}\bm{\sigma})Y_{t}+(\partial_{Y}\lambda_{t}^{*})\sigma_{Y}(\bm{\sigma}\bm{\rho})\right\}.$
This expression involves the costate $\lambda_{t}^{*}$ and its derivatives $\partial_{x}\lambda_{t}^{*}$ and $\partial_{Y}\lambda_{t}^{*}$. The first term within the braces represents the myopic demand driven by the risk premium $(\bm{\alpha}\bm{\sigma})Y_{t}$, and the second term captures the intertemporal hedging demand arising from the state variable $Y_{t}$. If $n=1$, this simplifies to $\pi_{t}^{*}=-\frac{1}{\sigma^{2}X_{t}^{*}\partial_{x}\lambda_{t}^{*}}\{\lambda_{t}^{*}\alpha\sigma Y_{t}+\sigma(\partial_{Y}\lambda_{t}^{*})\sigma_{Y}\rho\}$. For the specific case without intermediate consumption ($C_{t}=0$), an analytic solution involving Riccati ODEs can be derived using the HJB framework (Kim & Omberg, , 1996), as detailed in Appendix B.1.
@@ 2.3.2 A Multi-Asset, Multi-Factor Model
where $\kappa_{Y}\in\mathbb{R}^{k\times k}$ is a mean-reversion matrix, $\bm{\theta}_{Y}\in\mathbb{R}^{k}$ is the long-run mean vector, and $\bm{\sigma}_{Y}\in\mathbb{R}^{k\times k}$ governs the factor volatilities and correlations (e.g., $\mathrm{Cov}(d\mathbf{W}_{t}^{Y})=\Phi dt$).
@@
Here, $\mathbf{W}_{t}^{X}\in\mathbb{R}^{n}$ is a Brownian motion with covariance matrix $\Psi dt$. The asset shocks $\mathbf{W}_{t}^{X}$ and factor shocks $\mathbf{W}_{t}^{Y}$ can be correlated, with $\mathrm{Cov}(dW_{t}^{X,i},dW_{t}^{Y,j})=\rho_{ij}dt$. Let $\rho\in\mathbb{R}^{n\times k}$ be the matrix of cross-correlations.
@@
$\bm{\pi}_{t}^{*}=-\;\frac{1}{\,X_{t}^{*}\,\bigl{(}\partial_{x}\lambda_{t}^{*}\bigr{)}\!}\;\Sigma^{-1}\;\Bigl{\{}\,\underbrace{\lambda_{t}^{*}\;\mathrm{diag}(\bm{\sigma})\,\mathbf{A}\,\mathbf{Y}_{t}}_{\text{Myopic Term}}~{}+~{}\underbrace{\mathrm{diag}(\bm{\sigma})\,\rho\,\bm{\sigma}_{Y}\bigl{(}\partial_{\mathbf{Y}}\lambda_{t}^{*}\bigr{)}^{\top}}_{\text{Hedging Term}}\Bigr{\}},$
where $\bm{\sigma}=(\sigma_{1},\dots,\sigma_{n})^{\top}$, $\Sigma=\mathrm{diag}(\bm{\sigma})\Psi\mathrm{diag}(\bm{\sigma})$ is the $n\times n$ asset return covariance, $\mathbf{A}=(\bm{\alpha}_{1},\dots,\bm{\alpha}_{n})^{\top}$ is the $n\times k$ matrix of factor loadings, and $\partial_{\mathbf{Y}}\lambda_{t}^{*}$ is the $k\times 1$ gradient vector of the costate with respect to the factors $\mathbf{Y}_{t}$. This explicitly separates the demand into a myopic part driven by the current expected excess returns ($\mathrm{diag}(\bm{\sigma})\,\mathbf{A}\,\mathbf{Y}_{t}$) and a hedging part driven by the sensitivity to factor changes ($\partial_{\mathbf{Y}}\lambda_{t}^{*}$) and their correlations ($\rho$, $\Psi$, $\Phi$).
@@ 3 A Gradient-Based Algorithm for Multi-Asset Policy Optimization
This dual output enables distinct algorithmic approaches detailed in the subsequent subsections.
@@ 3.1 Baseline PG-DPO for the Extended Objective
    Simulate Forward Paths: For each sampled initial state $i$, simulate a discrete-time trajectory $\{(X_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\}_{k=0}^{N}$ forward from $t_{k}=t_{0}^{(i)}$ to $t_{N}=T$, using the current policy networks $(\bm{\pi}_{\theta},C_{\phi})$ and correctly correlated Brownian increments $(\Delta\mathbf{W}_{k}^{(i),X},\Delta\mathbf{W}_{k}^{(i),Y})$ derived from the Cholesky factor $L$ of $\mathbf{\Sigma}_{\text{block}}$. The state evolution follows:
@@
    where $\Delta t^{(i)}=(T-t_{0}^{(i)})/N$.
@@
    $J^{(i)}(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)};\theta,\phi)=\sum_{k=0}^{N-1}e^{-\delta t_{k}^{(i)}}U\bigl{(}C_{\phi}(t_{k}^{(i)},X_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\bigr{)}\Delta t^{(i)}+\kappa e^{-\delta T}U(X_{N}^{(i)}).$
@@
2:Policy nets $\pi_{\theta},\,C_{\phi}$; learning‑rate schedule $\{\alpha_{j}\}_{j=0}^{J_{\max}-1}$; max iters $J_{\max}$, batch size $M$, time steps $N$; initial‑state distribution $\eta$; utility $U$, discount $\delta$, terminal weight $\kappa$; market params $(\bm{\mu},\bm{\sigma},r,\bm{\mu}_{Y},\bm{\sigma}_{Y},\bm{\Omega})$
@@
18:         $J^{(i)}=\sum_{k=0}^{N-1}e^{-\delta t_{k}}U(C_{k})\Delta t+\kappa e^{-\delta T}U(X_{N})$
@@
Applying the chain rule one step backward ($\lambda_{k}^{(i)}=\lambda_{k+1}^{(i)}\frac{\partial X_{k+1}^{(i)}}{\partial X_{k}^{(i)}}$) and using a first-order expansion yields the following backward recurrence relation along the path $(i)$:
@@
This equation inherently possesses the structure of a discrete BSDE operating along the specific simulation path $(i)$. To formally connect this to the BSDE formulation in PMP, which involves processes adapted to the filtration $\{\mathcal{F}_{t_{k}}\}$ (representing available information at time $t_{k}$), one defines the adapted costate $\lambda_{k}=\mathbb{E}[\lambda_{k}^{(i)}|\mathcal{F}_{t_{k}}]$ and the predictable martingale component $Z_{k}$ (related to the conditional expectation of $\lambda_{k+1}^{(i)}\Delta W_{k}^{(i)}$). Taking the conditional expectation ($\mathbb{E}[\cdot|\mathcal{F}_{t_{k}}]$) of the pathwise recurrence relation above, and carefully defining $Z_{k}$ using a covariation-based approach, leads precisely to the standard discrete BSDE for the adapted process $\lambda_{k}$:
@@
Nevertheless, the fact that BPTT provides access to these theoretically grounded estimates is fundamental. Firstly, they are essential inputs for the Two-Stage PG-DPO variant discussed in the next subsection (§3.2). The baseline algorithm thus lays the groundwork by generating these estimates. Secondly, this deep connection explains why the algorithm is termed Pontryagin-Guided. Even though the baseline update only uses $\nabla_{(\theta,\phi)}J^{(i)}(\dots)$, the calculation of this gradient via BPTT structurally mirrors the backward propagation of sensitivities inherent in PMP’s adjoint equations (BSDEs), using the costate estimates $\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$ implicitly within the chain rule (cf. Huh, , 2024). Consequently, the optimization process, driven by these BPTT-computed gradients, is implicitly guided by the optimality conditions derived from Pontryagin’s principle, aiming to maximize the expected objective $\widetilde{J}$.
3.2 Two-Stage PG-DPO Algorithm
While the baseline PG-DPO algorithm described in Section 3.1 effectively optimizes the extended objective $\widetilde{J}$ by training the policy networks $(\bm{\pi}_{\theta},C_{\phi})$ end-to-end, it can sometimes require a significant number of iterations for the networks to fully converge. The Two-Stage PG-DPO variant offers a potentially more computationally efficient alternative by exploiting an empirical observation: the costate estimates ($\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$ and their relevant derivatives with respect to states, like $\partial_{x}\lambda_{k}^{(i)},\partial_{\mathbf{Y}}\lambda_{k}^{(i)}$), which are generated during BPTT, often stabilize and become reasonably accurate much faster than the full policy networks achieve optimality.
Algorithm 2 Two‑Stage PG‑DPO: Warm‑up Training & Deployment

### v3

@@ 2.2 Optimality Conditions via Pontryagin’s Maximum Principle
with initial conditions $X_{0}^{*}=x_{0}>x_{\min}$ and $\mathbf{Y}_{0}=\mathbf{y}_{0}\in\mathbb{R}^{k}$. Recall the definitions of the volatility matrix $\bm{\sigma}(t,\mathbf{Y}_{t})$ (for assets) and the covariance structure $\bm{\Omega}$ of the combined Brownian motion $\mathbf{W}_{t}=(\mathbf{W}^{X\top}_{t},\mathbf{W}^{Y\top}_{t})^{\top}$.
@@
The application of PMP involves maximizing a Hamiltonian function. To properly account for the dependence of the wealth diffusion on the control $\bm{\pi}$, the Hamiltonian should incorporate the interaction between the costate dynamics (represented by $\mathbf{Z}$ terms) and the diffusion coefficients. Following formulations in stochastic control theory (e.g., Yong and Zhou, , 2012), the appropriate Hamiltonian $\mathcal{H}$ can be expressed as:
$\mathcal{H}\!=\!e^{-\delta t}U(C)+\lambda\!\left[x\bigl{(}r+\pi^{\top}(\mu-r\mathbf{1})\bigr{)}-C\right]+\bm{\eta}^{\top}\mu_{Y}+\mathbf{Z}^{\lambda X}\!\bigl{(}x\sigma^{\top}\pi\bigr{)}+\bigl{(}\mathbf{Z}^{\eta X}\bigr{)}^{\!\top}\!\bigl{(}x\sigma^{\top}\pi\bigr{)}+\bigl{(}\mathbf{Z}^{\eta Y}\bigr{)}^{\!\top}\sigma_{Y}.$ (8)
For the $\bm{\pi}$-FOC, only the drift term $\lambda x\bm{\pi}^{\top}(\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1})$ and the diffusion-coupling term $\mathbf{Z}^{\lambda X}(x\,\bm{\sigma}(t,\mathbf{Y}_{t})^{\top}\bm{\pi})$ matter for the maximization step.
@@
    $C^{1}$-regular coefficients w.r.t. $\mathbf{y}$. All coefficient functions $\bm{\mu}_{Y},\,\bm{\sigma}_{Y},\,\bm{\mu},\,\bm{\sigma}\,$ $r$ are continuously differentiable with respect to the state vector $\mathbf{y}$. Furthermore, their partial derivatives ($\partial_{\mathbf{y}}f$) are globally Lipschitz continuous in $\mathbf{y}$ and satisfy linear growth conditions. This ensures that the drift terms in the costate BSDEs, involving derivatives like $\mathcal{H}_{\mathbf{y}}$, satisfy the required regularity conditions for the existence and uniqueness of solutions (cf. Yong and Zhou, , 2012).
@@
$\qquad\qquad-\frac{1}{X_{t}V_{xx}}\,\Sigma^{-1}(t,\mathbf{Y}_{t})\,\bigl{(}\bm{\sigma}(t,\mathbf{Y}_{t})\,\bm{\rho}\,\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})\bigr{)}V_{x\mathbf{y}}.$
Here, $\Sigma(t,\mathbf{Y}_{t})=\bm{\sigma}(t,\mathbf{Y}_{t})\Psi\bm{\sigma}(t,\mathbf{Y}_{t})^{\top}$ represents the instantaneous covariance matrix of the risky asset returns. The value function derivatives $V_{x}=\lambda_{t}^{*}$, $V_{xx}=\partial_{x}\lambda_{t}^{*}$, and $V_{x\mathbf{y}}=\nabla_{\mathbf{y}}\lambda_{t}^{*}$ (a $k\times 1$ vector) are evaluated along the optimal path $(t,X_{t}^{*},\mathbf{Y}_{t})$. The formula clearly separates the portfolio into two components: the myopic demand, which optimizes the immediate risk-return trade-off as if the investment opportunity set were constant, and the intertemporal hedging demand, which adjusts the portfolio to mitigate risks arising from future stochastic changes in the state variables $\mathbf{Y}_{t}$. The hedging demand’s magnitude and direction depend on the correlation $\rho$ between asset returns and state variable innovations, and crucially on $V_{x\mathbf{y}}$, which measures how the marginal value of wealth changes with the state variables.
@@
$\pi_{t}^{*}=-\frac{V_{x}}{X_{t}V_{xx}\sigma^{2}}\bigl{(}\mu(t,Y_{t})-r(t,Y_{t})\bigr{)}-\frac{1}{X_{t}V_{xx}\sigma^{2}}(\sigma\rho\sigma_{Y})V_{xy},$
@@ 2.3.1 A Multi-Asset Model with One OU Factor
We first consider a model with $n$ risky assets whose risk premia are driven by a single Ornstein–Uhlenbeck (OU) factor. Specifically, each asset $i$ has a constant volatility $\sigma_{i}>0$ and a time-varying risk premium proportional to an OU state variable $Y_{t}$, such that the expected return is $r+\alpha_{i}\sigma_{i}Y_{t}$, where $r$ is the constant risk-free rate. The state variable $Y_{t}$ follows the OU process:
@@
Let $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{n})^{\top}$ be the vector of loading coefficients and $\bm{\sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{n})$ be the diagonal matrix of constant asset volatilities, where each $\sigma_{i}>0$. The Brownian motions driving the assets, $\mathbf{W}_{t}^{S}=(W_{1,t}^{S},\dots,W_{n,t}^{S})^{\top}$, have an $n\times n$ covariance matrix $\Psi dt$, i.e., $\mathrm{Cov}(d\mathbf{W}_{t}^{S})=\Psi\,dt$. Furthermore, each $dW_{i,t}^{S}$ may have a correlation $\rho_{i}$ with $dW_{t}^{Y}$; let $\bm{\rho}=(\rho_{1},\dots,\rho_{n})^{\top}$.
@@
$dX_{t}=\bigl{[}X_{t}\bigl{(}r+\bm{\pi}_{t}^{\!\top}Y_{t}(\bm{\sigma}\bm{\alpha})\bigr{)}-C_{t}\bigr{]}dt+X_{t}\,\bm{\pi}_{t}^{\!\top}\bm{\sigma}\,d\mathbf{W}_{t}^{S}.$
The instantaneous covariance matrix of the $n$-asset returns is $\Sigma=\bm{\sigma}\,\Psi\,\bm{\sigma}$.
@@
$\bm{\pi}_{t}^{*}=-\frac{1}{X_{t}^{*}(\partial_{x}\lambda_{t}^{*})}\Sigma^{-1}\left\{\lambda_{t}^{*}Y_{t}(\bm{\sigma}\bm{\alpha})+(\partial_{Y}\lambda_{t}^{*})\sigma_{Y}(\bm{\sigma}\bm{\rho})\right\}.$
This expression involves the costate $\lambda_{t}^{*}$ and its derivatives $\partial_{x}\lambda_{t}^{*}$ and $\partial_{Y}\lambda_{t}^{*}$. The first term within the braces represents the myopic demand driven by the risk premium $Y_{t}(\bm{\sigma}\bm{\alpha})$, and the second term captures the intertemporal hedging demand arising from the state variable $Y_{t}$. If $n=1$, this simplifies to $\pi_{t}^{*}=-\frac{1}{\sigma^{2}X_{t}^{*}\partial_{x}\lambda_{t}^{*}}\{\lambda_{t}^{*}\alpha\sigma Y_{t}+\sigma(\partial_{Y}\lambda_{t}^{*})\sigma_{Y}\rho\}$. For the specific case without intermediate consumption ($C_{t}=0$), an analytic solution involving Riccati ODEs can be derived using the HJB framework (Kim and Omberg, , 1996), as detailed in Appendix B.1.
@@ 2.3.2 A Multi-Asset, Multi-Factor Model
where $\kappa_{Y}\in\mathbb{R}^{k\times k}$ is a mean-reversion matrix, $\bm{\theta}_{Y}\in\mathbb{R}^{k}$ is the long-run mean vector, $\bm{\sigma}_{Y}$ is the $k\times k$ diagonal matrix of factor-specific diffusion coefficients (as defined in Section 2.1), and $d\mathbf{W}_{t}^{Y}$ is a $k$-dimensional vector of Brownian motions whose components may be correlated, with $\mathrm{Cov}(d\mathbf{W}_{t}^{Y})=\Phi dt$.
@@
Here, $\mathbf{W}_{t}^{X}\in\mathbb{R}^{n}$ is a Brownian motion with covariance matrix $\Psi dt$. The asset shocks $\mathbf{W}_{t}^{X}$ and factor shocks $\mathbf{W}_{t}^{Y}$ can be correlated, with $\mathrm{Cov}(dW_{t}^{X,i},dW_{t}^{Y,j})=\rho_{ij}dt$. Let $\bm{\rho}\in\mathbb{R}^{n\times k}$ be the matrix of cross-correlations.
@@
$\bm{\pi}_{t}^{*}=-\;\frac{1}{\,X_{t}^{*}\,\bigl{(}\partial_{x}\lambda_{t}^{*}\bigr{)}\!}\;\Sigma^{-1}\;\Bigl{\{}\,\lambda_{t}^{*}\;\bm{\sigma}\,\mathbf{A}\,\mathbf{Y}_{t}~{}+~{}\bm{\sigma}\,\bm{\rho}\,\bm{\sigma}_{Y}\partial_{\mathbf{Y}}\lambda_{t}^{*}\Bigr{\}},$
where $\bm{\sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{n})$ is the $n\times n$ diagonal matrix of asset volatilities, $\Sigma=\bm{\sigma}\Psi\bm{\sigma}$ is the $n\times n$ asset return covariance, $\mathbf{A}=(\bm{\alpha}_{1},\dots,\bm{\alpha}_{n})^{\top}$ is the $n\times k$ matrix of factor loadings (where each $\bm{\alpha}_{i}$ is a $k\times 1$ vector of loadings for asset $i$), $\bm{\sigma}_{Y}$ is the $k\times k$ diagonal matrix of factor diffusion coefficients, and $\partial_{\mathbf{Y}}\lambda_{t}^{*}$ is the $k\times 1$ gradient vector of the costate with respect to the factors $\mathbf{Y}_{t}$. This explicitly separates the demand into a myopic part driven by the current expected excess returns ($\bm{\sigma}\,\mathbf{A}\,\mathbf{Y}_{t}$) and a hedging part driven by the sensitivity to factor changes ($\partial_{\mathbf{Y}}\lambda_{t}^{*}$) and their correlations ($\bm{\rho}$, $\Psi$, $\Phi$).
@@ 3 A Gradient-Based Algorithm for Multi-Asset Policy Optimization
This dual output enables distinct algorithmic approaches: the baseline PG-DPO method is detailed in Section 3.1, and the advanced “Two-Stage PG-DPO” variant, which leverages these costate estimates more directly, is explored in Section 3.2.
@@ 3.1 Baseline PG-DPO for the Extended Objective
    Simulate Forward Paths: For each sampled initial state $i$, simulate a discrete-time trajectory $\{(X_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\}_{k=0}^{N}$ forward from $t_{k}=t_{0}^{(i)}$ to $t_{N}=T$, using the current policy networks $(\bm{\pi}_{\theta},C_{\phi})$. The state evolution follows:
@@
    where $\Delta t^{(i)}=(T-t_{0}^{(i)})/N$, $\bm{\sigma}_{Y}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})$ is the vector of diffusion coefficients applied element-wise to the components of $\Delta\mathbf{W}_{k}^{(i),Y}$, and the correctly correlated Brownian increments $(\Delta\mathbf{W}_{k}^{(i),X},\Delta\mathbf{W}_{k}^{(i),Y})$ are derived from the Cholesky factor $L$ of the overall covariance matrix $\bm{\Omega}$.
@@
    $J^{(i)}(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)};\theta,\phi)=\sum_{k=0}^{N-1}e^{-\delta(t_{k}^{(i)}-t_{0}^{(i)})}U\bigl{(}C_{\phi}(t_{k}^{(i)},X_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\bigr{)}\Delta t^{(i)}+\kappa e^{-\delta(T-t_{0}^{(i)})}U(X_{N}^{(i)}).$
@@
2:Policy nets $\pi_{\theta},\,C_{\phi}$; learning‑rate schedule $\{\alpha_{j}\}_{j=0}^{J_{\max}-1}$; max iters $J_{\max}$, batch size $M$, time steps $N$; initial‑state distribution $\eta$; utility $U$, discount $\delta$, terminal weight $\kappa$; market params $(\bm{\mu},\bm{\sigma},r,\bm{\mu}_{Y},\bm{\sigma}_{Y}\text{ (k-dim vector of diff. coeffs.)},\bm{\Omega})$
@@
18:         $J^{(i)}=\sum_{k=0}^{N-1}e^{-\delta(t_{k}-t_{0}^{(i)})}U(C_{k})\Delta t+\kappa e^{-\delta(T-t_{0}^{(i)})}U(X_{N})$
@@
Applying the chain rule one step backward ($\lambda_{k}^{(i)}=\lambda_{k+1}^{(i)}\frac{\partial X_{k+1}^{(i)}}{\partial X_{k}^{(i)}}$) and using a first-order expansion yields the following backward recurrence relation along the path $(i)$ (this illustrative recurrence primarily shows influences from the wealth dynamics; a full derivation of the drift for $\lambda_{k}^{(i)}$ from the total reward $J^{(i)}$ would also account for terms related to consumption $U(C_{k}^{(i)})$ and the discount factor in its definition within the Hamiltonian context):
@@
This equation inherently possesses the structure of a discrete BSDE operating along the specific simulation path $(i)$. To formally connect this to the BSDE formulation in PMP, which involves processes adapted to the filtration $\{\mathcal{F}_{t_{k}}\}$ (representing available information at time $t_{k}$), one defines the adapted costate $\lambda_{k}=\mathbb{E}[\lambda_{k}^{(i)}|\mathcal{F}_{t_{k}}]$ and the predictable martingale component $Z_{k}$ (related to the conditional expectation of $\lambda_{k+1}^{(i)}\Delta W_{k}^{(i)}$). Taking the conditional expectation ($\mathbb{E}[\cdot|\mathcal{F}_{t_{k}}]$) of a more complete pathwise recurrence relation (inclusive of all terms affecting $J^{(i)}$) and carefully defining $Z_{k}$ using a covariation-based approach, leads precisely to the standard discrete BSDE for the adapted process $\lambda_{k}$:
@@
Nevertheless, the fact that BPTT provides access to these theoretically grounded estimates is fundamental. Firstly, they are essential inputs for the Projected PGDPO variant discussed in the next subsection (§3.2). The baseline algorithm thus lays the groundwork by generating these estimates. Secondly, this deep connection explains why the algorithm is termed Pontryagin-Guided. Even though the baseline update only uses $\nabla_{(\theta,\phi)}J^{(i)}(\dots)$, the calculation of this gradient via BPTT structurally mirrors the backward propagation of sensitivities inherent in PMP’s adjoint equations (BSDEs), using the costate estimates $\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$ implicitly within the chain rule (cf. Huh, , 2024). Consequently, the optimization process, driven by these BPTT-computed gradients, is implicitly guided by the optimality conditions derived from Pontryagin’s principle, aiming to maximize the expected objective $\widetilde{J}$.
3.2 Projected PG-DPO (P-PGDPO) Algorithm
While the baseline PG-DPO algorithm described in Section 3.1 effectively optimizes the extended objective $\widetilde{J}$ by training the policy networks $(\bm{\pi}_{\theta},C_{\phi})$ end-to-end, it can sometimes require a significant number of iterations for the networks to fully converge. The Projected PG-DPO (P-PGDPO) variant offers a potentially more computationally efficient alternative by exploiting an empirical observation: the costate estimates ($\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$ and their relevant derivatives with respect to states, like $\partial_{x}\lambda_{k}^{(i)},\partial_{\mathbf{Y}}\lambda_{k}^{(i)}$), which are generated during BPTT, often stabilize and become reasonably accurate much faster than the full policy networks achieve optimality. This P-PGDPO approach leverages these rapidly stabilizing costate estimates to construct near-optimal controls directly by projecting them onto the manifold of policies satisfying the PMP first-order conditions.
Algorithm 2 Projected PG‑DPO (P-PGDPO): Warm‑up Training & Deployment